### v1

# Search for spatial coincidences between galaxy mergers and Fermi-LAT 4FGL-DR4 sources

Abstract
We present a systematic search for spatial association between a high-confidence sample of 3,166 morphologically selected galaxy mergers detected through SDSS imaging and the 4FGL-DR4 catalog of gamma-ray sources detected by the Fermi Large Area Telescope (LAT). Using a conservative 4 positional uncertainty threshold and a Poisson-based statistical framework, we identify 21 statistically significant associations with match probabilities . Among these are known classes of gamma-ray emitters such as flat-spectrum radio quasars (FSRQs), BL Lacertae objects (BLL), and radio galaxies bolstering the hypothesis that merger-driven processes can fuel high-energy activity. Intriguingly, five of the associated sources remain unclassified in 4FGL-DR4, hinting at a possible link between galaxy mergers and a hitherto unrecognized population of gamma-ray sources. The dominance of AGN-like matches supports the scenario in which mergers trigger accretion onto central supermassive black holes, initiating AGN activity observable at gamma-ray energies. Moreover, the recurrent presence of unassociated sources among secure matches underscores the potential of merger catalogs as physically motivated priors in gamma-ray source identification efforts. This work constitutes the first dedicated effort to explore associations between gamma-ray sources and a large, morphologically selected sample of galaxy mergers, opening new avenues for understanding the role of interactions in high-energy astrophysics. We additionally examined the sample of 70 galaxy pairs from the Canadian Network for Observational Cosmology Field Galaxy Redshift Survey and found no statistically significant matches, with within the 4 positional uncertainty threshold.
I Introduction
The Fermi Large Area Telescope (LAT) has been a cornerstone in the exploration of the gamma-ray universe at GeV energies since its launch in 2008 Atwood et al. (2009) 111https://fermi.gsfc.nasa.gov/science/instruments/lat.html. It has provided an unprecedented view of the high-energy sky, enabling the detection and characterization of thousands of gamma-ray sources. The most recent data release, the Fourth Fermi-LAT Source Catalog Data Release 4 (4FGL-DR4) provides a comprehensive catalog of 7,195 gamma-ray sources observed over 14 years from 50 MeV to 1 TeV Ballet et al. (2023) (see also the related 4FGL-DR3 catalog Abdollahi et al. (2022)). These sources include a variety of astrophysical objects, such as blazars, radio galaxies, pulsars. However, a large number of Fermi-LAT sources () remain unidentified and no counterparts at lower energies have been identified. Numerous studies have therefore searched for spatial associations of Fermi 4FGL-DR4 sources with multi-wavelength source catalogs, including X-ray catalogs from SRG/eROSITA Mayer and Becker (2024) and SWIFT Ulgiati et al. (2025), radio source catalogs Bruzewski et al. (2023), SPT point sources at millimeter wavelengths Zhang et al. (2022), magnetars Ramakrishnan and Desai (2024), IceCube neutrino events Negro et al. (2023), MeerKAT sources Himes et al. (2025) etc.
In this work, we search for statistically significant association between Fermi 4FGL-DR4 catalog and galaxy mergers. Galaxy mergers are pivotal events in the evolution of galaxies, driving significant morphological and dynamical changes that influence star formation, black hole growth, and active galactic nuclei (AGN) activity Hopkins et al. (2006); Kaviraj (2025). These processes can produce energetic phenomena at gamma-ray energies, which can be detected by LAT. Galaxy mergers are thought to trigger intense star formation and fuel supermassive black holes, potentially leading to relativistic jets or other mechanisms responsible for gamma-ray production. Understanding the potential connection between galaxy mergers and gamma-ray sources is crucial for unraveling the physical processes governing these extreme astrophysical environments. A systematic search for associations between a comprehensive, high-confidence catalog of galaxy mergers and the latest Fermi-LAT data release has not yet been done, to the best of our knowledge.
This work is heavily inspired by Bouri et al. (2025), which searched for a spatial association between IceCube-detected neutrinos and sources from six different galaxy merger catalogs, since neutrinos could be produced through shock acceleration from galaxy mergers, as discussed in various works Lisenfeld and Völk (2010); Kashiyama and Mészáros (2014); Yuan et al. (2018). Along with the associated neutrinos, one would also expect high-energy gamma-ray radiation, which could be detected by Fermi-LAT. We briefly review some of the gamma-ray production mechanisms discussed in literature. The first estimate of gamma-ray flux from galaxy merger candidates was done in Lisenfeld and Völk (2010). This work considered two specific merger systems, viz. UGC 12914/5 and UGC 813/6 Lisenfeld and Völk (2010). These candidates are located at distances of 60 Mpc and 69 Mpc, respectively. The hadronic gamma-ray flux from decays resulting from energetic proton collisions with gas nuclei in the shocked region has been estimated in this fork to be , with a spectral index of 1.1 normalized at 1 TeV Lisenfeld and Völk (2010). Although this flux is too small to be detected by Fermi-LAT, for a galaxy merger happening at distances between 2.6 and 3.9 Mpc, this flux could be up to 700 times higher, thereby enhancing the detection probability of mergers in current generation gamma-ray telescopes such as Fermi-LAT. Then, Kashiyama and Mészáros (2014) investigated the shock acceleration of particles in galaxy mergers with stellar masses of and showed that cosmic rays could be accelerated upto 0.1-1 EeV with a spectral index of 2.0. It was shown that for a cosmic ray acceleration efficiency of 10%, one could detect gamma-ray radiation from about 10 nearby merging galaxies with Cherenkov Telescope Array Kashiyama and Mészáros (2014). In addition to UGC 12914/5 and UGC 813/6, another galaxy merger candidate, viz. VV 114 located at a distance of Mpc Vorontsov-Velyaminov et al. (2001) was identified as a potential gamma ray source. Subsequently, Yuan et al. (2018) extended this analysis and calculated the cumulative diffuse gamma-ray background by incorporating mergers of galaxy cluster and group candidates, and also incorporated the redshift evolution of average galactic radius, shock velocity and gas content inside haloes and galactic magnetic fields. This model is also consistent with the non-blazar contribution to extragalactic gamma-ray background Yuan et al. (2018).
The empirical study of galaxy mergers and correctly identifying galaxies that have been merging has traditionally been quite challenging (See Kaviraj (2025) for a recent review). Recent advances in catalogs from large-scale optical surveys, such as the Sloan Digital Sky Survey Data Release 16 (SDSS DR16) Ahumada et al. (2020), have however enabled the identification of large samples of galaxy mergers through morphological and spectroscopic signatures. Using machine learning techniques and high-confidence classification thresholds, Ackermann et al. (2018) constructed a galaxy merger catalog. This catalog contains a high-confidence sample of 3,166 galaxy mergers derived from a catalog of 328,151 candidates in the redshift range Ackermann et al. (2018). The second catalog we considered for our analysis is the J/AJ/130/2043 catalog, from the Canadian Network for Observational Cosmology field galaxy redshift survey, contains 70 close galaxy pairs across four sky regions covering 1.5 square degrees, with redshifts in the range Patton et al. (2005). We use both these aformentioned catalogs for spatial coincidence with Fermi-LAT point sources. These catalogs were also used for spatial coincidence analysis with IceCube neutrinos in Bouri et al. (2025). We note that other major merger catalogs Hwang and Chang (2009); Gimeno et al. (2004) used in Bouri et al. (2025) lack essential photometric or distance information required for flux-based probability estimates. Furthermore, several prominent merger systems discussed in the literature as promising gamma-ray sources, including UGC 12914/5, UGC 813/6, VV 114, NGC 660, and NGC 3256, are not present in the galaxy merger catalogs used in this study Ackermann et al. (2018). We also searched for spatial associations of these systems with sources in the 4FGL-DR4 catalog within a radius but found no matches. A targeted analysis of such individual systems will be conducted in a follow-up study.
II Data Analysis
We analyze the aformentioned galaxy merger catalog that contains 328,151 candidate mergers Ackermann et al. (2018). Each galaxy image in the catalog was processed by a convolutional neural network (CNN) classifier, which outputs a merger probability score , representing the likelihood that the system is undergoing a merger (with indicating a non-interacting galaxy and indicating a merger). The classifier was trained using 4-fold cross-validation, producing four independent estimates of for each galaxy Ackermann et al. (2018). To isolate high-confidence systems, we applied a strict threshold of across all four classifier outputs, retaining only those galaxies that satisfied this criterion in every fold. This conservative approach yielded a subset of 3,711 robust merger candidates. This stricter criterion minimizes the inclusion of ambiguous cases and emphasizes classifier agreement, albeit at the cost of reducing the final sample size. These selected objects were then cross-matched with the SDSS DR16 catalog Ahumada et al. (2020) using TOPCAT version 4.10-3 Taylor (2005), employing a matching radius of 0.1 arcseconds. The resulting cross-match identified 3,185 mergers with SDSS counterparts. Among these matched sources, 1,147 have measured spectroscopic redshifts and 3,166 sources possess SDSS -band apparent magnitudes. To characterize the spatial extent of the merger sample, we estimated the total angular sky area over which the 3,166 galaxy mergers are distributed. This effective footprint was computed by dividing the sky into finer bins in right ascension and declination and summing up the area of all the spatial bins that contained at least one merger. The contribution of each bin was corrected for spherical projection by weighting with cos(), where is the central declination of the bin. This produced an estimated sky coverage of approximately , which was used for surface density estimates and association probability calculations. Figure 1 shows the sky distribution in equatorial coordinates using an Aitoff projection.
For the gamma-ray dataset, we utilized 4FGL-DR4 catalog, which contains 7,195 gamma-ray sources detected between 50 MeV and 1 TeV over 14 years of observations Ballet et al. (2023).
To identify possible associations, we searched for candidate galaxy mergers located within the 4 positional uncertainty ellipse of each 4FGL-DR4 source. This conservative threshold, similar to that used in Zhang et al. (2022), accounts for both Fermi-LAT’s positional uncertainties and uncertainties in the external catalogs. When multiple optical sources were found within a single gamma-ray error ellipse, we assessed the probability of random association using a Poisson-based framework as described below Browne and Cohen (1978); Downes et al. (1986); Biggs et al. (2011); Zhang et al. (2022).
The expected number of chance associations ) is given by Browne and Cohen (1978); Downes et al. (1986); Biggs et al. (2011); Zhang et al. (2022):
where and are the semi-major and semi-minor axes of the Fermi positional ellipse, and is the surface density of mergers with flux greater than that of the SDSS counterpart having flux equal to . To evaluate , we used the SDSS-derived fluxes computed from the -band magnitudes of the merger candidates. For each candidate to be tested, we determined the number of mergers in the full sample with -band fluxes exceeding the corresponding candidate flux value. The details of the calculation of -band flux from the estimated magnitude can be found in Appendix A. Dividing this count by the total sky area covered by the catalog yielded a flux-dependent surface density. This value of was then used in Eq. 1 to calculate the expected number of chance associations (), allowing us to assess the probability of random coincidence.
The probability of random coincidence (i.e., the p-value) is then given by:
A small value indicates a low probability of random association, which implies a more reliable counterpart match. We adopt a threshold of to define a “secure” association, with a secondary “relaxed” threshold of also examined, following the criteria in Zhang et al. (2022). Given the rarity and low surface density of galaxy mergers, we expect the false match rate to remain low even with a relaxed threshold Downes et al. (1986); Biggs et al. (2011).
III Results
We began our analysis by examining a supplementary sample of 70 galaxy pairs from the Canadian Network for Observational Cosmology Field Galaxy Redshift Survey, as defined in Patton et al. (2005). Applying the same matching criterion and statistical methodology, we found no statistically significant associations with Fermi 4FGL-DR4 sources within the 4 positional uncertainty threshold . Therefore, we do not find any Fermi 4GL-DR4 counterparts in the J/AJ/130/2043 merger catalog.
Next, we applied the matching criterion defined in Equation 2 to our primary sample of 3,166 morphologically selected galaxy mergers identified from SDSS imaging. This resulted in identification of 38 spatial coincidences between Fermi 4FGL-DR4 sources and galaxy mergers from our sample. To assess the likelihood of true physical associations versus chance alignments, we computed Poisson-based match probabilities for each pair using Eq. 2. Among these 38 matches, 21 exhibit p-values below 0.05, indicating a low probability of random association. We classify these as statistically secure matches. According to source classifications in the 4FGL-DR4 catalog, the 21 matches within the 4 ellipses include four Flat Spectrum Radio Quasars (FSRQs), six BL Lacertae objects (BLL), three Blazar Candidates of Uncertain type (BCUs), and three Radio Galaxies (RDGs). The remaining five are unassociated sources (UIDs) lacking confirmed multiwavelength counterparts. Notably, none of the secure matches is classified as a pulsar. The five UIDs span gamma-ray flux densities in the range , suggesting the possible existence of a previously unrecognized population of gamma-ray sources associated with galaxy mergers.
Overall, our analysis reveals that approximately 0.67% (21 out of 3,166) of galaxy mergers exhibit statistically robust associations with gamma-ray sources. The predominance of blazar-type objects among these secure matches accounting for 62% is consistent with their known dominance in the extragalactic gamma-ray sky. However, the presence of radio galaxies and a significant fraction of unassociated sources (24%) suggests that merger-driven gamma-ray activity may originate from a broader range of physical mechanisms than previously recognized. To visualize these associations, Figure 2 presents a multi-panel view of Fermi sources with at least one galaxy merger candidate within the corresponding positional uncertainty ellipse. In each panel, the ellipse denotes the Fermi positional uncertainty region, the blue dot marks the Fermi source position, and the red cross indicates the galaxy merger counterpart with the lowest p-value. In Table 1 we depict all the secure matches with their coordinates, SDSS -mag, spectroscopic redshifts, flux densities, p-values, angular separation and classification. To explore marginal cases, we extended the analysis to include matches with , identifying 24 potential associations. Of these, 21 are already included in the secure sample (). The remaining three additional sources are 4FGL J1454.0+4927 (BCU), 4FGL J1523.2+2818 (BCU), and 4FGL J1223.3+1213 (BLL), with gamma-ray flux densities of , , and , and corresponding p-values of 0.069, 0.071, and 0.076, respectively. These sources are spatially associated with the SDSS galaxies SDSS J145506.73+493614.7, SDSS J152228.74+281402.9, and SDSS J122256.10+121739.2, which have spectroscopic redshifts of 0.0945, 0.0715, and 0.0905, respectively.
IV Conclusions
We have presented the first dedicated cross-matching analysis between two galaxy merger catalogs and the Fermi-LAT 4FGL-DR4 gamma-ray source catalog Ballet et al. (2023). The first catalog consists of a sample of 70 galaxy pairs from the Canadian Network for Observational Cosmology Field Galaxy Redshift Survey and found no statistically significant associations within the 4 positional uncertainty threshold (). The second catalog comprises of a high-confidence sample of 3,166 morphologically selected galaxy mergers Ackermann et al. (2018) and the Fermi-LAT 4FGL-DR4 gamma-ray source catalog Ballet et al. (2023). Using a conservative 4 positional uncertainty criterion and a Poisson-based statistical framework, we identified 38 spatial coincidences, of which 21 exhibit p-values below 0.05. These matches predominantly include well-known gamma-ray source classes, FSRQs, BL Lacs, and radio galaxies, supporting the hypothesis that merger-driven processes can fuel AGN activity and high-energy emission. Notably, five secure associations are with unclassified (unassociated) Fermi sources from the fourth FGL-DR4 catalog, suggesting that merging systems may host previously unrecognized gamma-ray populations.
While blazar-like AGN dominate the matched sample, the presence of radio galaxies, UIDs and the broad range of observed gamma-ray flux densities point to a diversity of emission mechanisms. These may include merger-induced starbursts, large-scale shocks, or obscured AGN activity, highlighting the need for deeper multi-wavelength follow-up Torres (2004); Ackermann et al. (2012). The consistent appearance of unassociated sources within the secure sample underscores the value of galaxy merger catalogs as informative priors for gamma-ray source identification. As gamma-ray localization continues to improve and future optical surveys such as Vera Rubin LSST, Euclid, and DESI expand the census of galaxy mergers, the ability to resolve such associations will significantly increase. We shall also conduct a targeted search for some specific merger candidates, such as UGC 12914/5, UGC 813/6, and VV 114 in a forthcoming work.
Fermi Source | Merger RA | Merger Dec | -mag | z | Flux (erg/cm2/s/Hz) | -value | Separation | Class |
4FGL J0016.20016 | 3.9988 | 0.3049 | 16.498 | 0.0396 | 0.0081 | 2.50 | FSRQ | |
4FGL J0813.4+2531 | 122.8163 | 25.1794 | 14.146 | 0.0137 | 0.0113 | 3.99 | FSRQ | |
4FGL J0815.6+3641 | 123.8059 | 36.7766 | 16.851 | 0.0420 | 0.0252 | 3.77 | FSRQ | |
4FGL J0844.5+3035 | 131.0952 | 30.5990 | 15.716 | 0.0748 | 0.0042 | 1.65 | UID | |
4FGL J1008.0+0028 | 151.9620 | 0.5303 | 17.075 | 0.0934 | 0.0111 | 3.13 | BLL | |
4FGL J1018.9+1043 | 154.6662 | 10.7793 | 14.790 | 0.0513 | 0.0058 | 0.53 | BCU | |
4FGL J1105.8+3944 | 166.4743 | 39.7825 | 15.958 | 0.0991 | 0.0026 | 2.26 | UID | |
4FGL J1112.0+2607 | 168.0185 | 26.0447 | 17.046 | 0.0696 | 0.0182 | 2.64 | UID | |
4FGL J1116.6+2915 | 169.1460 | 29.2497 | 14.657 | 0.0500 | 0.0006 | 0.80 | RDG | |
4FGL J1144.9+1937 | 176.2744 | 19.6365 | 17.479 | 0.0162 | 0.0083 | 1.97 | RDG | |
4FGL J1145.50340 | 176.3724 | 3.6637 | 16.811 | 0.0790 | 0.0122 | 0.83 | BLL | |
4FGL J1202.9+5141 | 180.7901 | 51.6782 | 15.849 | 0.0607 | 0.0060 | 1.67 | BCU | |
4FGL J1245.8+0232 | 191.5785 | 2.6659 | 17.394 | 0.0811 | 0.0344 | 3.31 | BLL | |
4FGL J1353.2+3740 | 208.3187 | 37.7076 | 16.675 | 0.0630 | 0.0076 | 1.26 | BLL | |
4FGL J1411.8+5249 | 212.9420 | 52.7960 | 16.248 | 0.0766 | 0.0028 | 2.64 | BLL | |
4FGL J1435.4+3338 | 218.7278 | 33.8222 | 15.394 | 0.0586 | 0.0171 | 3.00 | UID | |
4FGL J1441.7+1836 | 220.5586 | 18.6907 | 15.233 | 0.0710 | 0.0035 | 3.91 | BCU | |
4FGL J1508.8+2708 | 227.1547 | 27.1589 | 15.109 | 0.0760 | 0.0005 | 3.41 | BLL | |
4FGL J1518.6+0614 | 229.5286 | 6.3552 | 16.892 | 0.0795 | 0.0274 | 2.68 | RDG | |
4FGL J1600.4+0407 | 240.2073 | 4.1848 | 17.213 | 0.0877 | 0.0137 | 3.42 | UID | |
4FGL J1606.5+2717 | 241.6173 | 27.1373 | 17.119 | 0.0458 | 0.0480 | 3.70 | FSRQ |
Acknowledgements.
SM gratefully acknowledges the Ministry of Education (MoE), Government of India, for their consistent support through the research fellowship, which has been instrumental in facilitating the successful completion of this work. The authors also thank Lizhong Zhang for useful correspondence.References
- Atwood et al. (2009) W. B. Atwood, A. A. Abdo, M. Ackermann, W. Althouse, B. Anderson, M. Axelsson, L. Baldini, J. Ballet, D. L. Band, G. Barbiellini, et al., Astrophys. J. 697, 1071 (2009), eprint 0902.1089.
- Ballet et al. (2023) J. Ballet, P. Bruel, T. H. Burnett, B. Lott, and The Fermi-LAT collaboration, arXiv e-prints arXiv:2307.12546 (2023), eprint 2307.12546.
- Abdollahi et al. (2022) S. Abdollahi, F. Acero, L. Baldini, J. Ballet, D. Bastieri, R. Bellazzini, B. Berenji, A. Berretta, E. Bissaldi, R. D. Blandford, et al., Astrophys. J. Suppl. Ser. 260, 53 (2022), eprint 2201.11184.
- Mayer and Becker (2024) M. G. F. Mayer and W. Becker, Astron. & Astrophys. 684, A208 (2024), eprint 2401.17295.
- Ulgiati et al. (2025) A. Ulgiati, S. Paiano, F. Pintore, T. D. Russell, B. Sbarufatti, C. Pinto, E. Ambrosi, A. D’Aì, G. Cusumano, and M. Del Santo, Astron. & Astrophys. 694, A176 (2025), eprint 2412.19314.
- Bruzewski et al. (2023) S. Bruzewski, F. K. Schinzel, and G. B. Taylor, Astrophys. J. 943, 51 (2023), eprint 2212.05115.
- Zhang et al. (2022) L. Zhang, J. D. Vieira, M. Ajello, M. A. Malkan, M. A. Archipley, J. Capota, A. Foster, and G. Madejski, Astrophys. J. 939, 117 (2022), eprint 2206.06228.
- Ramakrishnan and Desai (2024) V. Ramakrishnan and S. Desai, arXiv e-prints arXiv:2412.03900 (2024), eprint 2412.03900.
- Negro et al. (2023) M. Negro, M. Crnogorčević, E. Burns, E. Charles, L. Marcotulli, and R. Caputo, Astrophys. J. 951, 83 (2023), eprint 2304.10934.
- Himes et al. (2025) M. Himes, P. Jagannathan, D. A. Frail, F. Schinzel, N. Gupta, S. A. Balashev, F. Combes, P. P. Deka, H. R. Klöckner, E. Momjian, et al., Astrophys. J. 980, 81 (2025), eprint 2412.12470.
- Hopkins et al. (2006) P. F. Hopkins, L. Hernquist, T. J. Cox, T. Di Matteo, B. Robertson, and V. Springel, Astrophys. J. Suppl. Ser. 163, 1 (2006), eprint astro-ph/0506398.
- Kaviraj (2025) S. Kaviraj, arXiv e-prints arXiv:2506.09136 (2025), eprint 2506.09136.
- Bouri et al. (2025) S. Bouri, P. Parashari, M. Das, and R. Laha, Phys. Rev. D 111, 063059 (2025), eprint 2404.06539.
- Lisenfeld and Völk (2010) U. Lisenfeld and H. J. Völk, Astron. & Astrophys. 524, A27 (2010), eprint 1009.1659.
- Kashiyama and Mészáros (2014) K. Kashiyama and P. Mészáros, Astrophys. J. Lett. 790, L14 (2014), eprint 1405.3262.
- Yuan et al. (2018) C. Yuan, P. Mészáros, K. Murase, and D. Jeong, Astrophys. J. 857, 50 (2018), eprint 1712.09754.
- Vorontsov-Velyaminov et al. (2001) B. A. Vorontsov-Velyaminov, R. I. Noskova, and V. P. Arkhipova, Astronomical and Astrophysical Transactions 20, 717 (2001).
- Ahumada et al. (2020) R. Ahumada, C. Allende Prieto, A. Almeida, F. Anders, S. F. Anderson, B. H. Andrews, B. Anguiano, R. Arcodia, E. Armengaud, M. Aubert, et al., Astrophys. J. Suppl. Ser. 249, 3 (2020), eprint 1912.02905.
- Ackermann et al. (2018) S. Ackermann, K. Schawinski, C. Zhang, A. K. Weigel, and M. D. Turp, MNRAS 479, 415 (2018), eprint 1805.10289.
- Patton et al. (2005) D. R. Patton, J. K. Grant, L. Simard, C. J. Pritchet, R. G. Carlberg, and K. D. Borne, Astron. J. 130, 2043 (2005), eprint astro-ph/0507417.
- Hwang and Chang (2009) C.-Y. Hwang and M.-Y. Chang, Astrophys. J. Suppl. Ser. 181, 233 (2009).
- Gimeno et al. (2004) G. N. Gimeno, R. J. Díaz, and G. J. Carranza, Astron. J. 128, 62 (2004).
- Taylor (2005) M. B. Taylor, in Astronomical Data Analysis Software and Systems XIV, edited by P. Shopbell, M. Britton, and R. Ebert (2005), vol. 347 of Astronomical Society of the Pacific Conference Series, p. 29.
- Browne and Cohen (1978) I. W. A. Browne and A. M. Cohen, MNRAS 182, 181 (1978).
- Downes et al. (1986) A. J. B. Downes, J. A. Peacock, A. Savage, and D. R. Carrie, MNRAS 218, 31 (1986).
- Biggs et al. (2011) A. D. Biggs, R. J. Ivison, E. Ibar, J. L. Wardlow, H. Dannerbauer, I. Smail, F. Walter, A. Weiß, S. C. Chapman, K. E. K. Coppin, et al., MNRAS 413, 2314 (2011), eprint 1012.0305.
- Torres (2004) D. F. Torres, Astrophys. J. 617, 966 (2004), eprint astro-ph/0407240.
- Ackermann et al. (2012) M. Ackermann, M. Ajello, A. Allafort, L. Baldini, J. Ballet, D. Bastieri, K. Bechtol, R. Bellazzini, B. Berenji, E. D. Bloom, et al., Astrophys. J. 755, 164 (2012), eprint 1206.1346.
- Lupton et al. (1999) R. H. Lupton, J. E. Gunn, and A. S. Szalay, Astron. J. 118, 1406 (1999), eprint astro-ph/9903081.
- Stoughton et al. (2002) C. Stoughton, R. H. Lupton, M. Bernardi, M. R. Blanton, S. Burles, F. J. Castander, A. J. Connolly, D. J. Eisenstein, J. A. Frieman, G. S. Hennessy, et al., Astron. J. 123, 485 (2002).
Appendix A SDSS Flux Conversion Methodology
SDSS magnitudes are reported in the asinh magnitude system Lupton et al. (1999), which handle low S/N and negative fluxes more robustly than traditional Pogson magnitudes. The inverse asinh formula for converting -band apparent magnitudes () to fluxes is:
where is the reference flux (1 maggy by definition), and is the softening parameter for the -band (see Table 21 in Stoughton et al. (2002)).
For completeness, the forward asinh magnitude definition is:
The resulting fluxes, in units of nanomaggies, were converted to Janskys using the SDSS relation:
and subsequently to cgs units via: